# Attention Retrieves, MLP Memorizes: Disentangling Trainable Components in the Transformer

### v1

@@ Abstract
The Transformer architecture is central to the success of modern Large Language Models (LLMs), in part due to its surprising ability to perform a wide range of algorithmic tasks – including mathematical reasoning, memorization, and retrieval – using only gradient-based training on next-token prediction. While the core component of a Transformer is the self-attention mechanism, we question how much, and which aspects, of the performance gains can be attributed to it. To this end, we compare standard Transformers to variants in which either the multi-layer perceptron (MLP) layers or the attention projectors (queries and keys) are frozen at initialization. To further isolate the contribution of attention, we introduce MixiT – the Mixing Transformer – a simplified, principled model in which the attention coefficients are entirely random and fixed at initialization, eliminating any input-dependent computation or learning in attention. Surprisingly, we find that MixiT matches the performance of fully trained Transformers on various algorithmic tasks, especially those involving basic arithmetic or focusing heavily on memorization. For retrieval-based tasks, we observe that having input-dependent attention coefficients is consistently beneficial, while MixiT underperforms. We attribute this failure to its inability to form specialized circuits such as induction heads – a specific circuit known to be crucial for learning and exploiting repeating patterns in input sequences. Even more interestingly, we find that attention with frozen key and query projectors is not only able to form induction heads, but can also perform competitively on language modeling. Our results underscore the importance of architectural heterogeneity, where distinct components contribute complementary inductive biases crucial for solving different classes of tasks. [^1] Our code is publicly available at https://github.com/princeton-pli/MixiT.
@@ 1 Introduction
  • We find that, in all but information retrieval and language modeling, MixiT achieves performance comparable to the fully trained transformer and Frozen-QK. These results suggest that for a wide range of tasks, "attention mixing is all you need" regardless of its specific form, in the sense that learned query-key interactions are not required.
  • When tested on an information retrieval task, MixiT fails to achieve good performance, while both the fully trained transformer and Frozen-QK can solve it to close-to perfect accuracy. We show that while the alternatives do form induction heads, MixiT cannot do so by design, and this strongly correlates with the observed (in)ability to solve the task.
  • On pure memorization tasks, Frozen-MLP loses two-thirds of the storage capacity of the fully-trainable Transformer, while Frozen-QK and MixiT only a third. This suggests that the MLP layers are largely responsible for memorization, and input-dependent attention has a very little added contribution.
@@ 3.1 Tasks
We benchmark on a variety of tasks, covering categories including basic mathematical reasoning, memorization, sentiment classification, and language modeling. The basic mathematical reasoning and memorization tasks are based on tasks used in [57], with increased difficulty on some tasks to better reflect differences between architectures.
@@
Each training instance encodes a small randomly generated sequence of pairs followed by a single query, and the model is required to emit the value associated with that query one step later. We sample a sequence length $m\sim\mathcal{U}\{1,\dots,m_{\max}\}$, where $m_{\max}$ is the maximum sequence length. We then sample $m$ keys $\{k_{i}\}_{i=1}^{m}$ iid from the set $\{\frac{V}{2},\dots,V\}$ and $m$ values $\{v_{i}\}_{i=1}^{m}$ from $\{1,\dots,\tfrac{\text{V}}{2}-1\}$, where $V=256$ is the vocabulary size. The resulting keys are interleaved with their values to form the prefix $[(k_{1},\,v_{1}),\,(k_{2},\,v_{2}),\,\dots,\,(k_{m},\,v_{m})]$. A query key $k_{q}$ with $q\in[m]$ is chosen uniformly at random from the keys, and appended to the sequence. The goal is to predict the value corresponding to the query token. This task isolates retrieval ability, and probes associative recall. We sample 40,000 sequences for training, and 4,000 sequences for testing. Transformer-based models typically solve this task by forming induction heads [36].
@@ 4 Results
Table 1 contains the main results on algorithmic and sentiment classification tasks with a large memorization focus.
@@
Trainable attention weights are necessary for retrieval.
We perform an experiment using the needle in a haystack dataset where the degree of complexity is controlled by the maximal number of pairs $m_{\max}$. We fix the embedding dimension $n=1024$ divided across $4$ heads with $n_{h}=256$, and fix 2 layers, which is sufficient to solve this task by using induction heads. We search for the optimal learning rate in the range $[5\cdot 10^{-5},5\cdot 10^{-3}]$. The results are shown in Figure 2. Notice how the while models with trainable queries and keys can solve the task across all the levels of difficulty, Frozen-QK has a drop in performances at $m_{\max}=40$ and MixiT struggles significantly after at $m_{\max}\geq 20$. These results provide compelling evidence that the MLP and embedding layers alone are insufficient when it comes to complex retrieval tasks.
Input-dependent attention patterns are required for language modeling.
@@
Memorization.
@@ 7 Conclusion
To complete the proof, we will invoke the Markov chain convergence to SDE results in the Skorohod topology, see for example Li et al. [28, Proposition A.6], which gives us the desired result.

### v2

@@ Abstract
The Transformer architecture is central to the success of modern Large Language Models (LLMs), in part due to its surprising ability to perform a wide range of algorithmic tasks – including mathematical reasoning, memorization, and retrieval – using only gradient-based training on next-token prediction. While the core component of a Transformer is the self-attention mechanism, we question how much, and which aspects, of the performance gains can be attributed to it. To this end, we compare standard Transformers to variants in which either the multi-layer perceptron (MLP) layers or the attention projectors (queries and keys) are frozen at initialization. To further isolate the contribution of attention, we introduce MixiT – the Mixing Transformer – a simplified, principled model in which the attention scores are entirely random and fixed at initialization, eliminating any input-dependent computation or learning in attention. Surprisingly, we find that MixiT matches the performance of fully trained Transformers on various algorithmic tasks, especially those involving basic arithmetic or focusing heavily on memorization. For retrieval-based tasks, we observe that having input-dependent attention scores is consistently beneficial, while MixiT underperforms. We attribute this failure to its inability to form specialized circuits such as induction heads – a specific circuit known to be crucial for learning and exploiting repeating patterns in input sequences. Even more interestingly, we find that attention with frozen key and query projectors is not only able to form induction heads, but can also perform competitively on language modeling. Our results underscore the importance of architectural heterogeneity, where distinct components contribute complementary inductive biases crucial for solving different classes of tasks. [^1] Our code is publicly available at https://github.com/princeton-pli/MixiT.
@@ 1 Introduction
  • We find that, in all but induction heads tasks and language modeling, MixiT achieves performance comparable to the fully trained transformer and Frozen-QK. These results suggest that for a wide range of tasks, "attention mixing is all you need" regardless of its specific form, in the sense that learned query-key interactions are not required.
  • When tested on tasks that require forming induction heads, MixiT fails to achieve good performance, while both the fully trained transformer and Frozen-QK can solve them to close-to perfect accuracy.
  • On pure memorization tasks, Frozen-MLP loses two-thirds of the storage capacity of the fully-trainable Transformer, while Frozen-QK and MixiT only a third. This suggests that the MLP layers are largely responsible, but collaborates with attention, on memorization tasks.
@@ 3.1 Tasks
We benchmark on a variety of tasks, covering categories including basic mathematical reasoning, memorization, sentiment classification, and language modeling. The basic mathematical reasoning and memorization tasks are based on tasks used in [59], with increased difficulty on some tasks to better reflect differences between architectures. More experimental details on these tasks can be found in Appendix B.
@@
Each training instance encodes a small randomly generated sequence of pairs followed by a single query, and the model is required to emit the value associated with that query one step later. We uniformly sample a sequence length $m\sim\mathcal{U}\{1,\dots,m_{\max}\}$, where $m_{\max}$ is the maximum sequence length. We then sample $m$ keys $\{k_{i}\}_{i=1}^{m}$ iid from the set $\{\frac{V}{2},\dots,V\}$ and $m$ values $\{v_{i}\}_{i=1}^{m}$ from $\{1,\dots,\tfrac{\text{V}}{2}-1\}$, where $V=256$ is the vocabulary size. The resulting keys are interleaved with their values to form the prefix $[(k_{1},\,v_{1}),\,(k_{2},\,v_{2}),\,\dots,\,(k_{m},\,v_{m})]$. A query key $k_{q}$ with $q\in[m]$ is chosen uniformly at random from the keys, and appended to the sequence. The goal is to predict the value corresponding to the query token. This task isolates retrieval ability, and probes associative recall. We sample 40,000 sequences for training, and 4,000 for testing. Transformer-based models typically solve this task by forming induction heads [37].
$k$-hop Induction Heads.
Following [42], the $k$-hop induction heads task, or hopk, recursively completes bigrams auto-regressively, by repeatedly predicting the token that followed the last occurrence of the currently-considered token. As an example: given the input $X=a{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}dc}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}ad}a$, the 2-hop induction heads prediction is ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}c}$. [42] showed that hopk is solvable by a $O(\log k)$-depth transformer. Hence to achieve fair comparison across architectures, we fix the model depth at 5 layers, and search over other hyperparameters.
@@ 4 Results
Table 1 contains the main results on algorithmic and sentiment classification tasks, many with a large memorization focus. Both Frozen-QK and MixiT are able solve such tasks. They are competitive with, and can even outperform, the standard fully trained transformer. These results highlight that input-dependent attention is not required for solving such algorithmic tasks.
@@
Frozen-QK and Frozen-MLP can solve induction heads tasks such as retrieval and hopk.
On the other side of the memorization-heavy tasks on the task spectrum, we test the model variations on tasks that require forming induction heads [17]: needle-in-a-haystack retrieval [37] and $k$-hop induction heads [42]. As shown in Table 2, Frozen-QK is able to solve the tasks on par with the standard fully trained transformer model, reinforcing the observation that induction heads can form even with frozen query and key projectors. On the other hand, MixiT is unable to solve these tasks, due to its inability to form specialized circuits as the attention scores are frozen at random initialization.
More detailed results on the retrieval task are shown in Figure 2, where we test the performance of each architecture at varying task complexities, controlled by the maximum number of key-value pairs $m_{\max}$ in the sequence. We search for the optimal learning rate for each architecture. To achieve a fair comparison across architectures, we fix the embedding dimension at $n=1024$, with $4$ heads and $2$ layers, which suffice to learn induction heads [17, 16]. As shown, Frozen-QK, Frozen-MLP, and the standard transformer all perform notably better than MixiT as the retrieval complexity increases; and Frozen-QK eventually deteriorates in quality faster than Frozen-MLP and the standard transformer, underscoring the key role attention plays in forming induction heads.
Task \ Model Standard Frozen-MLP Frozen-QK MixiT
$k$-hop Induction Heads $\uparrow$ 99.99% 99.89% 96.73% 48.58%
Retrieval $\uparrow$ 100% 100% 97.01% 11.24%
Table 2: Accuracies for tasks that require forming induction heads. Frozen-QK is able to solve the tasks on par with the standard fully trained transformer model, reinforcing the observation that induction heads can form even with frozen query and key projectors, e.g. as evident in Figure 3. Furthermore, as the attention module plays the key role in forming induction heads [17, 12], Frozen-MLP, with its trainable attention modules, is able to perform to almost perfect accuracy. On the other hand, MixiT is unable to solve these tasks, due to its inability to form specialized circuits as the attention scores are frozen at random initialization. The retrieval results are for sequences of maximum number of key-value pairs $m_{\text{max}}=30$.
Frozen-QK can perform competitively on language modeling.
@@
MLPs are crucial, and collaborate with attention, on memorization.
@@ 7 Conclusion
Acknowledgments
We are very grateful for the insightful discussions with Danqi Chen and Boris Hanin.
@@ Acknowledgments
To complete the proof, we will invoke the Markov chain convergence to SDE results in the Skorohod topology, see for example Li et al. [29, Proposition A.6], which gives us the desired result.
@@ B.1 Hyperparameter selection
For the $k$-hop induction heads task, following [42], we generate data using a sequence length of 100, and a maximum of $k=16$ hops. We use a training set of 100,000 samples, a test set of 100 samples, and train for 5,000 steps.